Discuss the symmetry of the graph y = x^3 - 3x
Answer by stanbon(75887)

Discuss the symmetry of the graph
f(x) = x^3 - 3x
f(-x) = -x^3 + 3x
-f(-x) = x^3 - 3x
----
Since f(x) = -f(-x), the graph will be symmetric
to the origin.
-------------
Cheers,
Stan H.
----------